OpenStudy (anonymous):

The length of a diagonal of a square is 2.6 inches. Find the perimeter of the square. 8.45 inches 2.68 inches 7.36 inches 6.82 inches

OpenStudy (campbell_st):

ok... so draw a diagram  find the value of x using pythagoras's theorem. then multiply the answer by 4 to get the perimeter

OpenStudy (anonymous):

how am I supposed to use Pythagoras theorem without knowing a or b ? ):

OpenStudy (anonymous):

Because you know c!

OpenStudy (campbell_st):

well its a square, they both have the same value... so looking at the diagram you can say \[x^2 + x^2 = 2.6^2\]

OpenStudy (campbell_st):

all you need to do is 1. solve for x then multiply the answer by 4

OpenStudy (campbell_st):

umm when I did it I found \[2x^2 = 6.76\] which means \[x = \sqrt{3.38}\] then multiply that by 4

OpenStudy (anonymous):

7.36 (:

OpenStudy (campbell_st):

yep, thats it